# Computing Large 2D/3D Convex Regions of Obstacle-Free Space for Robotic Planning

### v2

@@ VII-C Comparison of Solving MVIE
For MVIE (11), we propose two algorithms, namely the SOCP-reformulation algorithm and the randomized algorithm specialized for 2-D case, which are presented in Sec. V and VI, respectively. The randomized algorithm yields an analytical solution, here we abbreviate it as RAN. Then to distinguish from the methods to be compared, we denote the proposed SOCP-reformulation algorithm which is solved by affine scaling method as CAS. We compare them with three methods: 1) The optimization method based on SDP formulation of MVIE in IRIS [1]. 2) The example  [^4] https://docs.mosek.com/latest/cxxfusion/examples-list.html#doc-example-file-lownerjohn-ellipsoid-cc of the cutting-edge solver Mosek [23] to computes the Lowner-John inner ellipsoidal approximations of a polytope. 3) A strategy of solving the SOCP form (35) of the MVIE by Mosek. As the example in Mosek formulate MVIE into a mixed conic quadratic and semidefinite problem, we denote it as Mosek SDP and denote another strategy using Mosek as Mosek SOCP. We use the default parameters for IRIS and Mosek.
@@
TABLE VI: Comparison of Precision $\psi_{\mathcal{E}}$ between Different Methods for Solving MVIE
@@
where ${A_{\mathcal{E}}}^{*}$, ${D_{\mathcal{E}}}^{*}$ and ${b_{\mathcal{E}}}^{*}$ are the coefficients of the maximum inscribed ellipsoid solved by each method, $A_{\mathcal{P}}$ and $b_{\mathcal{P}}$ are the parameters that define the halfspaces of the input polytope $\mathcal{P}$ defined in (10), and $\mathcal{L}_{\psi}(\cdot)$ is the function defined in (61). This precision $\psi_{\mathcal{E}}$ represents the proximity of the ellipsoids obtained by different methods to the most active halfspace constraint, which are constructed in the form of second-order cone constraint as (27b). The results are summarized in Fig. 10 and Tab. VI. By comparing the performance of Mosek SDP and Mosek SOCP, we can get the conclusion that transforming the commonly used SDP formulation to the SOCP formulation, which is presented in Sec. V-A, leads to a significant improvement in computational efficiency without sacrificing precision. Additionally, in the low-dimension massive-constraint case faced in this paper, the use of the affine scaling method avoids the requirement of solving a large-scale system of linear equations at each iteration, compared to the primal-dual interior point method used in Mosek. Therefore as the results demonstrate, CAS is capable of solving MVIE in less time while maintaining a comparable precision. Moreover, for the 2-D case, our proposed linear-time complexity algorithm RAN further enhances the computational efficiency by orders of magnitude and compute analytical results.
@@ VIII-A Dense Corridor for 2-D Whole-body Planning
We adopt a differential driven AgileX SCOUT MINI  [^5] https://global.agilex.ai/chassis/11 as the physical platform. The platform is equipped with an Intel NUC  [^6] https://simplynuc.com/product/nuc11phki7c-full/ with an Intel Core i7-1165G7 as onboard processor and a LiDAR sensor  [^7] https://ouster.com/products/hardware/os1-lidar-sensor integrated with an IMU for perception. We use FAST-LIO2 [59] for both pre-mapping the environment and real-time localization during execution of the planned trajectory. That is, the obstacles are represented by points.
@@
VIII-A1 Narrow Maze
@@ VIII-B Sparse Corridor for 3-D Local Replanning
Our experimental platform is a customized quadrotor equipped with an NVIDIA Orin NX [^8] https://www.nvidia.com/en-us/autonomous-machines as the onboard processor and a Livox MID-360999https://www.livoxtech.com/mid-360 LiDAR for perception. For the quadrotor, due to its small size, we model it as a point mass, similar to many other works [2, 8]. We employ FAST-LIO2 [59] for online state estimation and utilize occupancy grid map to filter sensor noise, which is easy to inflate for safety margins. For trajectory planning, the path generation and corridor generation are the same as Sec. VII-A3. Then building upon our previous work [3], we utilize piecewise polynomials to represent the trajectory. Each piece of the trajectory is constrained within the corresponding polytope to ensure safety.
@@ IX Discussion and Conclusion
Our algorithm FIRI aims to generate the largest possible convex polytopes. However, when it comes to practical applications such as trajectory planning, we cannot guarantee that the inflation direction of the polytope will always be favorable for trajectory optimization. For example, when the robot has a high velocity, if the inflation direction is perpendicular to the velocity direction, it may lead to the generated convex polytope that are not conducive to the trajectory optimization. Our future research direction will focus on generating convex polytopes that are application-friendly.
@@
In CGAL [60], there has been a lack of a dedicated solver for 2-D MVIE. In the future, we will continue our research on 2-D MVIE and work towards implementing a version of rational predicates.

### v3

@@ VII-C Comparison of Solving MVIE
For MVIE (11), we propose two algorithms, namely the SOCP-reformulation algorithm and the randomized algorithm specialized for 2-D case. The randomized algorithm yields an analytical solution, here we abbreviate it as RAN. To distinguish from above methods, we denote the proposed SOCP-reformulation algorithm which is solved by affine scaling method as CAS. We compare them with three methods: 1) The optimization method based on SDP formulation of MVIE in IRIS [1]. 2) The example  [^5] https://docs.mosek.com/latest/cxxfusion/examples-list.html#doc-example-file-lownerjohn-ellipsoid-cc of the cutting-edge solver Mosek [23] to computes the Lowner-John inner ellipsoidal approximations of a polytope. 3) A strategy of solving the SOCP form (35) of the MVIE by Mosek. As the example in Mosek formulate MVIE into a mixed conic quadratic and semidefinite problem, we denote it as Mosek SDP and denote another strategy using Mosek as Mosek SOCP. We use the default parameters for IRIS and Mosek.
where ${A_{\mathcal{E}}}^{*}$, ${D_{\mathcal{E}}}^{*}$ and ${b_{\mathcal{E}}}^{*}$ are the coefficients of the maximum inscribed ellipsoid solved by each method, $A_{\mathcal{P}}$ and $b_{\mathcal{P}}$ are the parameters that define the halfspaces of the input polytope $\mathcal{P}$ defined in (10), and $\mathcal{L}_{\psi}(\cdot)$ is the function defined in (61). This precision $\psi_{\mathcal{E}}$ represents the proximity of the ellipsoids obtained by different methods to the most active halfspace constraint, which are constructed in the form of second-order cone constraint as (27b). The results are summarized in Fig. 10 and Tab. V. By comparing the performance of Mosek SDP and Mosek SOCP, we can get the conclusion that transforming the commonly used SDP formulation to the SOCP formulation, which is presented in Sec. V-A, leads to a significant improvement in computational efficiency without sacrificing precision. Additionally, in the low-dimension massive-constraint case faced in this paper, the use of the affine scaling method avoids the requirement of solving a large-scale system of linear equations at each iteration, compared to the primal-dual interior point method used in Mosek. Therefore as the results demonstrate, CAS is capable of solving MVIE in less time while maintaining a comparable precision. Moreover, for the 2-D case, our proposed linear-time complexity algorithm RAN further enhances the computational efficiency by orders of magnitude and compute analytical results.
@@
TABLE V: Comparison of Precision $\psi_{\mathcal{E}}$ between Different Methods for Solving MVIE
@@ VIII-A Dense Corridor for 2-D Whole-body Planning
We adopt a differential driven AgileX SCOUT MINI  [^6] https://global.agilex.ai/products/scout-mini as the physical platform. The platform is equipped with an Intel NUC  [^7] https://www.intel.com/content/dam/support/us/en/documents/intel-nuc/NUC11PH_TechProdSpec.pdf with an Intel Core i7-1165G7 as onboard processor and a LiDAR sensor  [^8] https://ouster.com/products/hardware/os1-lidar-sensor integrated with an IMU for perception. We use FAST-LIO2 [64] for both pre-mapping the environment and real-time localization during execution of the planned trajectory. That is, the obstacles are represented by points.
@@
VIII-A1 Cluttered Environment
VIII-A2 Narrow Maze
@@ VIII-B Sparse Corridor for 3-D Local Replanning
Our platform is a customized quadrotor equipped with an NVIDIA Orin NX [^9] https://www.nvidia.com/en-us/autonomous-machines/ as the onboard processor and a Livox MID-360101010https://www.livoxtech.com/mid-360 LiDAR for perception. For the quadrotor, due to its small size, we model it as a point mass, similar to many other works [2, 8]. We employ FAST-LIO2 [64] for online state estimation and utilize occupancy grid map to filter sensor noise, which is easy to inflate for safety margins. For trajectory planning, the path generation and corridor generation are the same as Sec. VII-A3. Then building upon our previous work [3], we utilize piecewise polynomials to represent the trajectory. Each piece of the trajectory is constrained within the corresponding polytope to ensure safety.
@@ IX Discussion and Conclusion
Our algorithm FIRI aims to generate the largest possible convex polytopes. However, when it comes to practical applications such as trajectory planning, we cannot guarantee that the inflation direction of polytope is always favorable for trajectory optimization. For example, when the robot has a high velocity, if the inflation direction is perpendicular to the velocity direction, it may lead to generated polytope that are not conducive to trajectory optimization. Our future research direction will focus on generating convex polytopes that are application-friendly. Additionally, the number of faces of the generated polytope significantly affects how fast the motion planning problem can be solved. As discussed in Sec. III.B1, we currently employ a widely used greedy method to reduce this number, and in the future, we will explicitly incorporate the face number as a further criterion of the region quality, while balancing it with the volume of the polytope.
@@
CGAL [65] lacks a dedicated solver for 2-D MVIE. In the future, we will continue our research on 2-D MVIE and work towards implementing a version of rational predicates.